Need assistance finding answer.
Moving to another question will save this response. Table 1 For the following question(s), refer to the table below. class interval cum f 139 137 130 115 cum % 100.00 98.56 93.53 82.73 67.63 41.01 22.30 8.63 4.32 1.44 relative % 90-99 80-89 70-79 1.44 5.04 10.79 15.11 26.62 18.71 13.67 4.32 2.88 1.44 15 21 37 50-59 40-49 30-39 20-29 10-19 0-9 31 12 19 Refer to Table 1. N equals a. 139 b. 137 c. 145 d. 135
Answer #1

N = Total number of observations

N = 139

N = sum(f) = 2 + 7 +15 + ......+ 4 + 2 = 139

N = 139

N equals 139